OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x − 1) −12? x −5 x −5 −5 x A number line with a closed circle on negative 5 and shading to the right.

OpenStudy (anonymous):

You forgot the inequality signs when you were copying/pasting it :)

OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x − 1) −12? x >or= −5 x <or =−5 −5 <or = x A number line with a closed circle on negative 5 and shading to the right